OpenStudy (calculusxy):

medal!!! Marcus studies two plant cells in biology class. One cell measures 8 x 10^-4 centimeters across. The other cell measures 2 x 10^-3 centimeters across. Marcus wants to know how many times greater in size the larger cell is. He compares the sizes of the cells using these steps: 1.) Divide the whole number factors: 8 / 2 = 4 2.) Divide the powers of 10: 10^-3 / 10^-4 = 10^1 = 10 3.) Multiply the products: 4x10=40 He concludes that the larger cell is 40 times greater in size than the smaller cell. Is Marcus' solution process and conclusion correct? Explain.

OpenStudy (calculusxy):

I got the answer `2.5 times greater`.

jhonyy9 (jhonyy9):

than the first step is 8/4 so than the second need being 10^-4/10^-3 not invers

OpenStudy (calculusxy):

i did \(\large \frac{2 \times 10^{-3}}{8 \times 10^{-4}}\)

OpenStudy (calculusxy):

which would be 0.25 \(\times 10^1\) and then move it one space to the left to get \(2.5 \times 10^0\) which is the same thing as `2.5 times greater`

OpenStudy (danjs):

right, 10^1 / 4

OpenStudy (calculusxy):

is my answer and my work is correct?

OpenStudy (danjs):

yeah that works

OpenStudy (danjs):

10/4 = 5/2

OpenStudy (calculusxy):

thanks :)

OpenStudy (anonymous):

je pense que ce correct.....

OpenStudy (calculusxy):

Je vous remercie :)

OpenStudy (anonymous):

de rien :d

OpenStudy (calculusxy):

Il est très bien

OpenStudy (calculusxy):

Merci d'être mon deuxième personne pour vérifier!

OpenStudy (calculusxy):

Bonne nuit

jhonyy9 (jhonyy9):

2*10^-3 = 20*10^-4 20*10^-4 --------- = 20/8 = 5/2 = 2,5 times greater sure 8*10^-4 bonne nuit

jhonyy9 (jhonyy9):

@calculusxy